OpenStudy (anonymous):

solve Sqrt x+3-3=x the sqrt is only over the x+3

OpenStudy (anonymous):

x+3=(x+3)sqr sqrroot(x+3)=x+3 assuming x+3=y sqrroot(y)=y y can only be either 1 or o solving for x y=x+3 0=x+3 x=-3 y=x+3 1=x+3 x=-2 therefore the solutions are x=-2 and x=-3

OpenStudy (anonymous):

@jo12345 Thank you so much!